OpenStudy (anonymous):

The roots of the equation x2 - 10x + 25 = 0 are: Answer a. 2 real b. 2 imaginary c. 1 real d. no solution

OpenStudy (anonymous):

\(x^2 - 10x + 25 = x^2 -(5+5) x +5\times 5 \implies x = 5 \) 1 real.

OpenStudy (precal):